OpenStudy (walgenbajr12):

How would I solve 1/2-2/3x=1/4?

OpenStudy (shadowlegendx):

\[\frac{ 1 }{ 2 } - \frac{ 2 }{ 3 } x = \frac{ 1 }{ 4 }\]

OpenStudy (shadowlegendx):

Is this what it looks like

OpenStudy (walgenbajr12):

yes

OpenStudy (shadowlegendx):

Do you know what the first step is?

OpenStudy (walgenbajr12):

I made the common factor 12 and multipled throughout

OpenStudy (shadowlegendx):

What did you get?

OpenStudy (walgenbajr12):

12-24+12x=12

OpenStudy (shadowlegendx):

no no no

OpenStudy (shadowlegendx):

\[12(\frac{ 1 }{ 2 } - \frac{ 2 }{ 3 } x = \frac{ 1 }{ 4 })\]

OpenStudy (shadowlegendx):

We get:

OpenStudy (shadowlegendx):

\[\frac{ 12 }{ 2 } - \frac{ 24 }{ 3 }x = \frac{ 12 }{ 4 }\]

OpenStudy (shadowlegendx):

Do you understand how I got those numbers?

OpenStudy (walgenbajr12):

Yeah but I thought that because the bottom all has the common factor of 12 they become canceled out

OpenStudy (shadowlegendx):

They do, but not in the way you're thinking

OpenStudy (otherworldly):

why don't u just d this 1/2 - 1/4 = 2/3x and go from there

OpenStudy (walgenbajr12):

okay!

OpenStudy (shadowlegendx):

Remember \[12 = \frac{ 12 }{ 1 }\] Just multiply across

OpenStudy (shadowlegendx):

@OtherWorldly He should understand it this way first

OpenStudy (otherworldly):

ok

OpenStudy (walgenbajr12):

so now what do I do?

OpenStudy (shadowlegendx):

Going back to what I have way up there, we continue \[6 - 8x = 3\]

OpenStudy (shadowlegendx):

We got that from this \[\frac{ 12 }{ 2 } - \frac{ 24 }{ 3 }x = \frac{ 12 }{ 4 }\]

OpenStudy (walgenbajr12):

how???

OpenStudy (otherworldly):

divided

OpenStudy (shadowlegendx):

What part do you not understand?

OpenStudy (walgenbajr12):

oh okay got it!!

OpenStudy (shadowlegendx):

Alright we continue \[6 - 8x = 3\] \[-8x = -3\]

OpenStudy (walgenbajr12):

so it would just be 3/8/

OpenStudy (shadowlegendx):

Correct

OpenStudy (shadowlegendx):

\[x = \frac{ -3 }{ -8 }\]

OpenStudy (walgenbajr12):

Thank you so much!! Can you help me with one more problem??

OpenStudy (shadowlegendx):

Simplified \[x = \frac{ 3 }{ 8 }\]

OpenStudy (shadowlegendx):

Sure

OpenStudy (walgenbajr12):

-4/7b + 7=11 I have a bunch like this I just need to remember how to solve one of them so I can work on the rest. I forgot how to solve for b if theres a fraction. I tried taking 7-11 which equals -4 and tried dividing -4 by -4/7 to get B but my online homework checker says its wrong

OpenStudy (shadowlegendx):

Does it look like this? \[\frac{ -4}{ 7 }b + 7 = 11\]

OpenStudy (walgenbajr12):

correct

OpenStudy (shadowlegendx):

\[7(\frac{ -4 }{ 7 } +7 = 11)\]

OpenStudy (shadowlegendx):

\[-4b + 49 = 77\]

OpenStudy (shadowlegendx):

\[-4b = 28\]

OpenStudy (shadowlegendx):

\[b = -7\]

OpenStudy (walgenbajr12):

why did you multiply by 7

OpenStudy (shadowlegendx):

Let me rewrite that one step, I forgot the b. \[7(\frac{ -4 }{ 7 }b + 7 = 11)\]

OpenStudy (shadowlegendx):

To eliminate the fraction

OpenStudy (walgenbajr12):

ohhhh okay! Thank you so so much! You were a really big help. I was starting to get really frustrated with myself!

OpenStudy (shadowlegendx):

tbh I did it the hard way. You could subtract 7 first, from 11. -4/7b = 4 Multiply by 7 -4b = 28 b = -7

OpenStudy (shadowlegendx):

Same result

OpenStudy (shadowlegendx):

I personally like to do it the harder way to keep my mind doing harder calculations

OpenStudy (shadowlegendx):

But do it either way

OpenStudy (otherworldly):

i was wondering y u were doing it the hard way XD

OpenStudy (shadowlegendx):

I knew you were :P

OpenStudy (shadowlegendx):

I have class, I'll be back online later today if you need further help. Have a nice day guys, hope you get your work done ^-^

OpenStudy (otherworldly):

k good luck in class bye